OpenStudy (anonymous):

Pleeeasee help D: ! 3. Which inequality can be used to solve the following problem? Mona rented a car for $175 per week plus $0.25 per mile. How far did Mona travel if she spent at most $350 A. 175 + 0.25m < 350 B. 175 + 0.25m > 350 C. 175 + 350 < 0.25m D. 350 + 0.25m < 175

OpenStudy (anonymous):

175 + 0.25m < 350

OpenStudy (anonymous):

I would have liked it to be <= 350, because at most implies it could also equal 350.

OpenStudy (anonymous):

5. The selling price of a certain CD is $9 more than the price the store paid the manufacturing company. If the selling price is $32, which equation can be used to find the price, p, the store paid the manufacturing company? A. P + 9 = 32 C. P + 32 = 9 B. P - 9 = 32 D. 9 - p = 32

OpenStudy (anonymous):

@GT

OpenStudy (anonymous):

Post it separately. People can get to find it later.

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

175+0.25<=350

OpenStudy (anonymous):

p+9=32

OpenStudy (anonymous):

6. Solve: 4x + 2 (x - 6) = 30 Step 1: 4x + 2x - 6 = 30 Step 2: 6x - 6 = 30 Step 3: 6x = 30 Step 4: x = 5 Which is the first incorrect step in the solution shown above? A. Step 1 C. Step 3 B. Step 2 D. Step 4

OpenStudy (anonymous):

@Taufique